OpenStudy (anonymous):

Write the vertex form for a parabola with the given characteristics. vertex ( 0, 0) directrix x = -15 Help Please!

OpenStudy (campbell_st):

well the distance from the vertex to the directrix is the focal length so a = 15 even though the drectrix is x = -15 the distance from the vertex is 15 units. ans since the directrix is below the vertex the parabola is concave up. the parabola will be in the form \[(x -h)^2 = 4a(y -k)\] (h, k) is the vertex and a is the focal length. in your question you know h = 0 and k = 0 so just substitute h, k and a to find the parabola

OpenStudy (anonymous):

so I'm looking for Y

OpenStudy (campbell_st):

nope.. h = 0 and k = 0 gives \[(x -0)^2 = 4a(y - 0)... or ... x^2 = 4ay\] now substitute a = 15 into the equation \[x^2 = 4 \times 15 \times y\] just simplify it